What a common term used to refer to the measure of central tendency otherwise known as the mean.?
average
What has the author P B written?
P. B. has written:
'A discourse of baptism: or an answer to Mr. Stubs's sermon ... to Mr. Dorrington's Vindication of baptism; and to a discourse of the late Mr. Nat. Taylor. By P. B. ..'
'Christianity revived, and judaism subverted'
i think you roll all 10 of these dice at once and it is similar to farkle
What is the probability of having a normal child?
The probability is zero! There is no such thing as "normal". Every child (and adult) has some unique characteristics and that makes them not normal - in that respect.
What are your expectation in your PE 1?
Since I have nothing to do with PE 1 I have absolutely no expectations in it.
How do you find the odd of picking a green marble 3 yellow 4 blue 6 green and 7 pink marbles?
The probability, in one draw, is 6/20 or 30%
What are 3 outcomes as a result of reconstruction?
The north got what it really wanted: The union was saved, and the federal properties in the south were back in the possession of the federal government. Freed blacks were abandoned and allowed to drift back into virtual slavery, because ending slavery was never the reason the north mounted military operations against the south.
It was not a healing time. It was a time in which the Republican-dominated Congress forced some harsh conditions upon the south. Andrew Johnson wanted to carry on with Lincoln's relatively benign policies toward the post-war south, but Congress was in a mood for vengeance, so they impeached Johnson on trumped up charges in an attempt to neutralize his influence.
Later, when Grant became president, he sent U.S. troops south to enforce the integration laws just as Eisenhower did almost one hundred years later, but the northern population and certainly the southern population were not yet ready for such federal action, so every president backed off of such action until Eisenhower tried it. This attempt to use force to integrate blacks into mainstream society is a large part of the reason Grant was considered our worst president for so many years. (His brother stealing money from the Bureau of Indian Affairs didn't help President Grant's reputation any.)
Back then, reconstruction was a reference to reconstructing the Union; as soon as all of the Confederate states had been re-admitted to the Union, people back then considered reconstruction completed. It wasn't until later that reconstruction started to take on connotations of social reconstruction and integration.
What is the genotypic ratio of the offspring from a cross of AAA x AAA?
The genotypic ratio of a cross of Aa and Aa is: one AA, one aa, and two Aa. Or 1:2:1
There is insufficient information for us to even begin to understand this question. Please edit the question to include more context or relevant information.
Probability is a concept so there is no point trying to tell it anything. There is nothing that is capable of listening, let alone understanding.
If, by tell, you mean "find the probability" then you need to provide more information about what the event is and what else you know about it.
There is insufficient information for us to even begin to understand this question. Please edit the question to include more context or relevant information.
Probability is a concept so there is no point trying to tell it anything. There is nothing that is capable of listening, let alone understanding.
If, by tell, you mean "find the probability" then you need to provide more information about what the event is and what else you know about it.
There is insufficient information for us to even begin to understand this question. Please edit the question to include more context or relevant information.
Probability is a concept so there is no point trying to tell it anything. There is nothing that is capable of listening, let alone understanding.
If, by tell, you mean "find the probability" then you need to provide more information about what the event is and what else you know about it.
There is insufficient information for us to even begin to understand this question. Please edit the question to include more context or relevant information.
Probability is a concept so there is no point trying to tell it anything. There is nothing that is capable of listening, let alone understanding.
If, by tell, you mean "find the probability" then you need to provide more information about what the event is and what else you know about it.
The probability is approx 0.09. This assumes that J and K are not prime numbers.
When you roll 3 dice what are the odds of rolling one six?
The probability of rolling exactly one six is equal to:
1/6 * 5/6 * 5/6, which is equal to 25/216, or 11.5740741 percent.
What is the lowest number you can get when two dice are rolled?
The lowest number will be obtained by getting lowest number on both of the dice if its a 6 sided dice (normal kind) then the lowest will be 2
How many simple events are there in the sample space of four puppies?
There are 16 simple events in the sample space of four puppies.
1-6
What is the probability of having a child with an attached earlobe?
Whether or not the earlobe is attached is a genetically inherited trait and so you would need to look at earlobes in your, and your partner's, families.
Why simple random sampling is the basis for all probability samplings?
It is not: the question is misguided.
Why does a researcher want to go from a normal distribution to a standard normal distributio?
A normal distribution simply enables you to convert your values, which are in some measurement unit, to normal deviates. Normal deviates (i.e. z-scores) allow you to use the table of normal values to compute probabilities under the normal curve.
How many ways can you arrange the letters oversimplification?
> 6.40237371 × 1015
Actually, since there are four i's and two o's, the number of distinct permutations of the letters in "oversimplification" is 18!/(4!2!) = 133,382,785,536,000.
Is identical twins are more closely related than tyhe fraternal twins?
Yes. Identical twins started out as one egg and one sperm cell and are genetically identical. Fraternal twins come from two eggs and two sperm cells and are no more closely related than any other sibling.
What is the area between z1.35 and z 2.24 and z0?
Assuming that these refer to the standard normal distribution, the areas are as follows:
z = 0 to 1.35: 0.411
z = 0 to 2.24: 0.487
z = 1.35 to 2.24: 0.076
Assuming that these refer to the standard normal distribution, the areas are as follows:
z = 0 to 1.35: 0.411
z = 0 to 2.24: 0.487
z = 1.35 to 2.24: 0.076
Assuming that these refer to the standard normal distribution, the areas are as follows:
z = 0 to 1.35: 0.411
z = 0 to 2.24: 0.487
z = 1.35 to 2.24: 0.076
Assuming that these refer to the standard normal distribution, the areas are as follows:
z = 0 to 1.35: 0.411
z = 0 to 2.24: 0.487
z = 1.35 to 2.24: 0.076
What is the probability of rolling 3 dies not having the same number?
The probabilty is 3 out of 6 or 3/6
One way to look at it is: It doesn't matter what you get with the first roll (say a 5) but now there is only a 1 in 6 chance of rolling a 5, and a 1 in 6 chance with the third roll. So if you multiply the probabilty of the first roll (which being certain has a score of 1.0) by the two other probabilities (1/ 6 and 1/6) the overall probability is 1 * 1/6 * 1/6 = 1/36